OpenStudy (anonymous):

Solve the equation by graphing x^2-5x-24=0

OpenStudy (anonymous):

i got x=-3 and x=8

OpenStudy (solomonzelman):

then you have to say \(\normalsize\color{red}{ \bf y=x^2-5x-24 }\) plug in x values and get the coordinates just like you would get when doing a liner equation. example: \(\normalsize\color{red}{ \bf y=x^2-5x-24 }\) \(\normalsize\color{red}{ \bf y=(1)^2-5(1)-24 }\) \(\normalsize\color{red}{ \bf y=1-5-24 }\) \(\normalsize\color{red}{ \bf y=-28 }\) So one point is (-28) I am going to tell you also that it is a U shaped parabola.

OpenStudy (solomonzelman):

I meant to say that the point is (1, -28)

OpenStudy (solomonzelman):

BUT< solving this by graphing is indeed stupid !!!!

OpenStudy (anonymous):

i just typed the equation into a graphing calculator and found where the points crossed the x-axis and those are the solutions

OpenStudy (solomonzelman):

yes, b/c in your case, y=0,

OpenStudy (anonymous):

which happened to be (-3,0) and (8,0)

zepdrix (zepdrix):

Here is a nice graphing utility if you need one :) Bookmark this website or something. https://www.desmos.com/calculator/9vs3bweyet

OpenStudy (solomonzelman):

you wrote not y=, but 0=

OpenStudy (anonymous):

i just did what i knew, solve by graphing. put in calculator and look on the x-axis for solutions of an equation.

OpenStudy (solomonzelman):

yes, and these solutions are called the y-intercepts.

OpenStudy (anonymous):

do you mean x-intercepts? because they intercept with the x-axis not the y-axis.